The Cellar Division of House Company makes a number of electrical components that it sells to other manufacturers in their region. Cellar currently makes and sells 15,000 units per year of a particular item, #EGX 400. Cellar’s variable cost of production is $23/unit; this item normally sells for $45 in the open market. Cellar’s annual fixed costs are $250,000. Cellar has an annual manufacturing capacity of 30,000 units.

Beyer Company has offered to purchase 8,500 units from Cellar for a price of $30/unit. Cellar has declined this one-time offer.

  1. Briefly summarize the Cellar Division’s likely reasons for rejecting the offer.
  2. Prepare calculations to evaluate this offer using the differential costing method we have used in class.
  3. Do you agree with Cellar’s rejection of the offer? Explain.
Answer #1

Answer:

Sales revenue if 8,500 units are sold in an open market @ $45/unit = 8,500*45 = $382,500

Sales revenue if 8,500 units are sold to Beyer company @ $30/unit = 8,500*30 = $255,000

Differential revenue = $127,500

Conclusion: Since, the Cellar Division can earn $127,500 more if they sell these 8,500 units in an open market when compared selling them to Beyer company; they have rejected this offer.

I agree with this decision because Cellar Division makes more revenue by selling these 8,500 units in an open market.

Note: Variable costs and fixed costs are not considered since they would be same in both the scenarios as there is no change in such costs.